# A Construction of non-degenerate $\mathbb{Z}_{2}$-harmonic functions on $\mathbb{R}^{n}$

### v1

@@ 4. Asymptotic behavior of the $\mathbb{Z}_{2}$-harmonic functions
In other words, $f_{\boldsymbol{h}}-C$ is asymptotic to a harmonic quadric of index $n-1$.
@@
Notice that $\frac{\mathrm{d}}{\mathrm{d}y}yS(y)=n\prod_{l=1}^{n-1}(y+p_{l})$ and henceforth integration by part yields to
@@
Similarly use Equation 2.7, we conclude
@@
If we let $t^{-1}\boldsymbol{h}=(t^{-1}h_{i})$, then the $\mathbb{Z}_{2}$-harmonic function $t^{-1}f_{t^{-1}\boldsymbol{h}}$ has the same quadric term as $f_{\boldsymbol{h}}$ in asymptotic expansion, but the constant term becomes $t^{-1}a_{0}$, which will go to 00 as $t\to\infty$. Moreover, in this case, the ellipsoid scales down to the origin.
To prove the Proposition 1.2, we will need to show that any multiple of $n-1$ positive parameters can be reached by $(a_{1},\cdots,a_{n-1})$. To see this, define a cone $K_{n}$ in $\mathbb{R}^{n-1}$ as follows
@@
The map $F_{n}$ is surjective. In other words, for every index $n-1$ harmonic quadric, there exists a non-degenerate $\mathbb{Z}_{2}$-harmonic polynomial that asymptotes to it.
@@
It is easy to see that $F_{n}=(F_{n,1},\cdots,F_{n,n-1})$ can be extended continuously to the closure $\overline{K_{n}}$. Indeeds, $\partial\overline{K_{n}}$ is a union of $\overline{K_{n-1}}$ and $F_{n}$ can be extended continuously to $F_{n-1}$ on $\overline{K_{n-1}}$. Let $P_{n}$ be the radial projection from $\overline{K_{n}}\setminus\{0\}$ to the intersection
@@
then $F_{n}$ is surjective is equivalent to $P_{n}\circ F_{n}:\Delta_{n}\to\Delta_{n}$ is surjective. This amounts to show that $P_{n}\circ F_{n}|_{\partial\Delta_{n}}$ is homotopic to $Id|_{\partial\Delta_{n}}$, where the homotopy is
@@
When $n=3$, surjectivity reduced to intermediate value theorem. On the other hand, when $n>3$, we prove by contradiction, and induction. Suppose $P_{n}\circ F_{n}$ is not surjective. By induction hypothesis, let $p$ be a point in the interior of $\Delta_{n-1}$, which does not lies in the image of $P_{n}\circ F_{n}$. Now let $\pi_{p}$ be the radial projection from $\Delta_{n}\setminus\{p\}$ to $\partial\Delta_{n}$. Therefore,
@@
It is very likely that the map $F_{n}$ is bijective. We leave it to the interested readers.
@@ 5. Relation to Donaldson’s Twistor construction
In the recent paper [3], Simon Donaldson discovered another construction of $\mathbb{Z}_{2}$-harmonic function using the twistor method when $n=3$. In this section, we demonstrate that our example coincides with Donaldson’s when $n=3$. In his construction, the $\mathbb{Z}_{2}$-harmonic function is represented by a contour integral in a twistor space, which can be written as
@@
The branching set of $f$ is an ellipse in $x_{1}x_{2}$-plane
@@
Suppose $x_{3}\geq 0$, we can choose a single-valued branch of $\tan^{-1}(\frac{w}{\sqrt{Q}})$. Although the function $\tan^{-1}(\frac{w}{\sqrt{Q}})$ will develop logarithm singularities at $w=\pm\sqrt{-1}\sqrt{Q}$, the integrand remains bounded.
The main ingredient to the proof is the uniqueness theorem established by Weifeng Sun [7].
@@
If $K$ is a compact smoothly embedded oriented codimension-$2$ submanifold in $\mathbb{R}^{n},n\geq 3$. Then there is a $1$-$1$ correspondence (up to a sign) between a $\mathbb{Z}_{2}$-harmonic function $f$ that branches along $K$ with polynomial growth and a harmonic polynomial $P$. The bijection is given by
@@
To show two constructions agree when $n=3$, it suffices to prove that the asymptotic behaviors of the $\mathbb{Z}_{2}$-harmonic functions at large distance agree. We now estimate Donaldson’s examples as in [3]. Decompose $S^{1}$ into $I_{1}\sqcup I_{2}$
@@
Changing the variable in the integrand yields to
@@ 6. Lawlor’s necks with small angle
In a slightly different vein, the differential $\mathrm{d}f_{\boldsymbol{h}}$ defines a $\mathbb{Z}_{2}$-harmonic one form on $\mathbb{R}^{n}$. Regard $T^{*}\mathbb{R}^{n}=(x_{1},y_{1}\frac{\partial}{\partial x_{1}},\cdots,x_{n},y_{n}\frac{\partial}{\partial x_{n}})$ as $\mathbb{C}^{n}$ such that $z_{i}=x_{i}+\sqrt{-1}y_{i}$ and let $(\mathbb{C}^{n},\omega,g,\Omega)$ be the flat Calabi-Yau structure, where $g$ is the Euclidean metric on $\mathbb{C}^{n}$ and
@@
Moreover, the symplectic structure $\omega=\mathrm{d}\lambda$ is exact with Liouville form $\lambda=-\sum_{i=1}^{n}y_{i}\mathrm{d}x_{i}$. A special Lagrangian submanifold is a real $n$-dimensional submanifold such that
@@
We call a special Lagrangian exact if $\lambda|_{L}$ is an exact form. In this section, we will discuss the relation between the multi-valued graph $\mathrm{d}f_{\boldsymbol{h}}$ and a family of exact special Lagrangians, which are known as Lawlor’s necks [10, 11].
@@
A Lawlor’s necks is a smoothing of the singular special Lagrangian $\Pi_{\boldsymbol{0}}\cup\Pi_{\boldsymbol{\phi}}$ in the case when
@@
For our purpose, we will consider a model acted by
@@
The Lawlor’s necks $L_{\boldsymbol{c}}$ is constructed as follows. Let
@@
Here, $\phi_{i}$ satisfy the relation in Equation 6.2. Let
@@
accordingly. It can be checked that $m_{k}^{\prime}(s)>0,1\leq k\leq n$. As a consequence, the projection is a $2$-$1$ map away from a codimension-$2$ ellipsoid
@@
In other words, $L_{\boldsymbol{c}}$ can be viewed as a multivalued graph from $\Pi_{\boldsymbol{0}}=\mathbb{R}^{n}$ to $\mathbb{C}^{n}=T^{*}\mathbb{R}^{n}$. In particular, if we view $-\lambda_{L}$ as a multivalued one form on $\Pi_{0}$, then the multivalued graph is defined by $-\lambda_{L}$ tautologically. Moreover, since $L_{\boldsymbol{c}}$ is diffeomorphic to $S^{n-1}\times\mathbb{R}$, which is simply connected. The Liouville form is exact
@@
Now consider the case when $\phi_{i},1\leq i\leq n-1$ are sufficiently small. Define a family of multiple, parametrized by a sufficiently large positive number $t$
@@
The angles $\tilde{\phi}_{i}$ do not obey the equation 6.2. Consequently, there is no uniform choice of Lagrangian angle $\hat{\theta}$ that simultaneously makes $\tilde{\Pi}_{\pm}$ into special Lagrangians. But as $t\to+\infty$, they satisfy
@@
Let $U\Subset\mathbb{R}^{n}\setminus E_{\boldsymbol{h}}$ be a simply connected open set, we can choose a single valued branch of $F_{\boldsymbol{c}_{t}}$. It following from the fact that $\mathrm{d}F_{\boldsymbol{c}_{t}}$ defines a special Lagrangian, $tF_{\boldsymbol{c}_{t}}$ is a solution to
@@
Let $\sigma_{k}(Hess(f))$ be the $k$-th symmetric polynomials of the eigenvalues of $Hess(F)$. Expand the above equation
@@
where $k=\lfloor\frac{n-1}{2}\rfloor$. We will show that $tF_{\boldsymbol{c}_{t}}$ converges to a smooth function $F$ on $U$, and the the equation will imply $F$ is harmonic. The uniqueness theorem will implies $F=f_{\boldsymbol{h}}$. Consequently, $tF_{\boldsymbol{c}_{t}}$ converges to $f_{\boldsymbol{h}}$ in $C_{loc}^{\infty}(\mathbb{R}^{n})$.
To see this, we will need to understand the behavior of the projection map when $t$ is large. Recall that
@@
Similarly, we define $Pr_{Y}$ as the projection of the $2$-valued graph to the cotangent direction
@@
in $C_{loc}^{\infty}(\mathbb{R})$ with error bounded by $O(t^{-2})$. To understand the behavior of $\psi_{i},\psi_{n}$, we will consider two cases. The first one is $|s|$ is bounded from below and above and the second one is $|s|$ is sufficiently small.
@@
in $C^{\infty}_{loc}(\mathbb{R}\times S^{n-1})$ respectively, with error bounded by $O(t^{-2})$.
We define two kinds of open sets in $\mathbb{R}^{n}\setminus E_{\boldsymbol{h}}$, that exhaust $\mathbb{R}^{n}\setminus E_{\boldsymbol{h}}$
@@
As we have shown before, $F$ is $\mathbb{Z}_{2}$-harmonic on $\mathbb{R}^{n}\setminus E_{\boldsymbol{h}}$. In particular, $|\nabla F|,|F|$ are bounded near $E_{\boldsymbol{h}}$. We can extend $|F|$ to $E_{\boldsymbol{h}}$ continuously by setting $|F|=0$ on $E_{\boldsymbol{h}}$. It remains to show $F=f_{\boldsymbol{h}}$. Because $\mathrm{d}F$ and $\mathrm{d}f_{\boldsymbol{h}}$ have the same asymptotic behavior at infinity, then we obtain

### v2

@@ 4. Asymptotic behavior of the $\mathbb{Z}_{2}$-harmonic functions
In other words, $f_{\boldsymbol{h}}-C$ is asymptotic to a harmonic quadric of index n1n-1.
@@
Notice that $\frac{\mathrm{d}}{\mathrm{d}y}yS(y)=n\prod_{l=1}^{n-1}(y+p_{l})$ and henceforth, integration by parts yields to
@@
Similarly, using Equation 2.7, we conclude
@@
If we let $t^{-1}\boldsymbol{h}=(t^{-1}h_{i})$, then the $\mathbb{Z}_{2}$-harmonic function $t^{-1}f_{t^{-1}\boldsymbol{h}}$ has the same quadric term as $f_{\boldsymbol{h}}$ in the asymptotic expansion, but the constant term becomes $t^{-1}a_{0}$, which will go to 0 as $t\to\infty$. Moreover, in this case, the ellipsoid scales down to the origin.
To prove the Proposition 1.2, we will need to show that any multiple of n1n-1 positive parameters can be reached by $(a_{1},\cdots,a_{n-1})$. To see this, define a cone $K_{n}$ in $\mathbb{R}^{n-1}$ as follows
@@
The map $F_{n}$ is surjective. In other words, for every index n1n-1 harmonic quadric, there exists a non-degenerate $\mathbb{Z}_{2}$-harmonic polynomial that asymptotes to it.
@@
It is easy to see that $F_{n}=(F_{n,1},\cdots,F_{n,n-1})$ can be extended continuously to the closure $\overline{K_{n}}$. Indeed, $\partial\overline{K_{n}}$ is a union of $\overline{K_{n-1}}$ and $F_{n}$ that can be continuously extended to $F_{n-1}$ on $\overline{K_{n-1}}$. Let $P_{n}$ be the radial projection from $\overline{K_{n}}\setminus\{0\}$ to the intersection
@@
then $F_{n}$ is surjective if and only if $P_{n}\circ F_{n}:\Delta_{n}\to\Delta_{n}$ is surjective. This amounts to showing that $P_{n}\circ F_{n}|_{\partial\Delta_{n}}$ is homotopic to $Id|_{\partial\Delta_{n}}$, where the homotopy is
@@
When $n=3$, surjectivity is reduced to the intermediate value theorem. On the other hand, when $n>3$, we prove by contradiction and induction. Suppose $P_{n}\circ F_{n}$ is not surjective. By the induction hypothesis, let $p$ be a point in the interior of $\Delta_{n-1}$, which does not lie in the image of $P_{n}\circ F_{n}$. Now let $\pi_{p}$ be the radial projection from $\Delta_{n}\setminus\{p\}$ to $\partial\Delta_{n}$. Therefore,
@@
The injectivity of the map $F_{n}$ was recently shown by Yuanbo Zhou through direct computation [20]. In fact, to verify the injectivity, it suffices to show that the Jacobian $DF_{n}$ is injective on $K_{n}$. Let $\Phi=a_{0}\prod_{j=1}^{n-1}h_{i}^{-1}$, Yuanbo Zhou observed that
$\frac{1}{2}a_{i}=\frac{\partial}{\partial h_{i}}\Phi.$
The injectivity of $DF_{n}$ then follows immediately from the fact that the Hessian for the integrand of $\Phi$ is strictly positive with respect to the variables $h_{1},\cdots,h_{n-1}$.
@@ 5. Relation to Donaldson’s Twistor construction
In the recent paper [4], Simon Donaldson discovered another construction of $\mathbb{Z}_{2}$-harmonic functions using the twistor method when $n=3$. In this section, we demonstrate that our example coincides with Donaldson’s when $n=3$. In his construction, the $\mathbb{Z}_{2}$-harmonic function is represented by a contour integral in twistor space, which can be written as
@@
The branching set of $f$ is an ellipse in the $x_{1}x_{2}$-plane
@@
Suppose $x_{3}\geq 0$, we can choose a single-valued branch of $\tan^{-1}(\frac{w}{\sqrt{Q}})$. Although the function $\tan^{-1}(\frac{w}{\sqrt{Q}})$ will develop logarithmic singularities at $w=\pm\sqrt{-1}\sqrt{Q}$, the integrand remains bounded.
The main ingredient for the proof is the uniqueness theorem established by Weifeng Sun [11].
@@
If $K$ is a compact smoothly embedded oriented codimension-$2$ submanifold in $\mathbb{R}^{n},n\geq 3$. Then there is a 11-11 correspondence (up to a sign) between a $\mathbb{Z}_{2}$-harmonic function $f$ that branches along $K$ with polynomial growth and a harmonic polynomial $P$. The bijection is given by
@@
To show that two constructions agree when $n=3$, it suffices to prove that the asymptotic behaviors of the $\mathbb{Z}_{2}$-harmonic functions at large distances agree. We now estimate Donaldson’s examples as in [4]. Decompose $S^{1}$ into $I_{1}\sqcup I_{2}$
@@
Changing the variable in the integrand yields
@@ 6. Lawlor’s necks with small angle
In a slightly different vein, the differential $\mathrm{d}f_{\boldsymbol{h}}$ defines a $\mathbb{Z}_{2}$-harmonic one form on $\mathbb{R}^{n}$. Regard $T^{*}\mathbb{R}^{n}=(x_{1},y_{1}\frac{\partial}{\partial x_{1}},\cdots,x_{n},y_{n}\frac{\partial}{\partial x_{n}})$ as $\mathbb{C}^{n}$ such that $z_{i}=x_{i}+\sqrt{-1}y_{i}$, and let $(\mathbb{C}^{n},\omega,g,\Omega)$ be the flat Calabi-Yau structure, where $g$ is the Euclidean metric on $\mathbb{C}^{n}$ and
@@
Moreover, the symplectic structure $\omega=\mathrm{d}\lambda$ is exact with the Liouville form $\lambda=-\sum_{i=1}^{n}y_{i}\mathrm{d}x_{i}$. A special Lagrangian submanifold is a real $n$-dimensional submanifold such that
@@
We call a special Lagrangian exact if $\lambda|_{L}$ is an exact form. In this section, we will discuss the relationship between the multi-valued graph $\mathrm{d}f_{\boldsymbol{h}}$ and a family of exact special Lagrangians, which are known as Lawlor’s necks [17, 18].
@@
A Lawlor’s neck is a smoothing of the singular special Lagrangian $\Pi_{\boldsymbol{0}}\cup\Pi_{\boldsymbol{\phi}}$ in the case when
@@
For our purposes, we will consider a model acted by
@@
The Lawlor’s necks $L_{\boldsymbol{c}}$ are constructed as follows. Let
@@
Here, $\phi_{i}$ satisfies the relation in Equation 6.2. Let
@@
accordingly. It can be checked that $m_{k}^{\prime}(s)>0,1\leq k\leq n$. As a consequence, the projection is a 22-11 map away from a codimension-$2$ ellipsoid
@@
In other words, $L_{\boldsymbol{c}}$ can be viewed as a multivalued graph from $\Pi_{\boldsymbol{0}}=\mathbb{R}^{n}$ to $\mathbb{C}^{n}=T^{*}\mathbb{R}^{n}$. In particular, if we view $\lambda_{L}$ as a multivalued one form on $\Pi_{0}$, then the multivalued graph is defined tautologically by $\lambda_{L}$. Moreover, since $L_{\boldsymbol{c}}$ is diffeomorphic to $S^{n-1}\times\mathbb{R}$, which is simply connected. The Liouville form is exact
@@
Now consider the case when $\phi_{i},1\leq i\leq n-1$ is sufficiently small. Define a family of multiple, parameterized by a sufficiently large positive number $t$
@@
The angles $\tilde{\phi}_{i}$ do not obey the equation 6.2. Consequently, there is no uniform choice of the Lagrangian angle $\hat{\theta}$ that simultaneously makes $\tilde{\Pi}_{\pm}$ into special Lagrangians. But as $t\to+\infty$, they satisfy
@@
Let $U\Subset\mathbb{R}^{n}\setminus E_{\boldsymbol{h}}$ be a simply connected open set, we can choose a single valued branch of $F_{\boldsymbol{c}_{t}}$. It follows from the fact that $\mathrm{d}F_{\boldsymbol{c}_{t}}$ defines a special Lagrangian, $tF_{\boldsymbol{c}_{t}}$ is a solution to
@@
Let $\sigma_{k}(Hess(f))$ be the $k$-th symmetric polynomial of the eigenvalues of $Hess(F)$. Expand the above equation
@@
where $k=\lfloor\frac{n-1}{2}\rfloor$. We will show that $tF_{\boldsymbol{c}_{t}}$ converges to a smooth function $F$ on $U$, and the equation will imply $F$ is harmonic. The uniqueness theorem will imply $F=f_{\boldsymbol{h}}$. Consequently, $tF_{\boldsymbol{c}_{t}}$ converges to $f_{\boldsymbol{h}}$ in $C_{loc}^{\infty}(\mathbb{R}^{n})$.
To see this, we need to understand the behavior of the projection map when $t$ is large. Recall that
@@
Similarly, we define $Pr_{Y}$ as the projection of the $2$-valued graph onto the cotangent direction
@@
in $C_{loc}^{\infty}(\mathbb{R})$ with an error bounded by $O(t^{-2})$. To understand the behavior of $\psi_{i},\psi_{n}$, we will consider two cases. The first one, $|s|$, is bounded from below and above, and the second one, $|s|$, is sufficiently small.
@@
in $C^{\infty}_{loc}(\mathbb{R}\times S^{n-1})$ respectively, with an error bounded by $O(t^{-2})$.
We define two kinds of open sets in $\mathbb{R}^{n}\setminus E_{\boldsymbol{h}}$, which exhaust $\mathbb{R}^{n}\setminus E_{\boldsymbol{h}}$
@@
As we have shown before, $F$ is $\mathbb{Z}_{2}$-harmonic on $\mathbb{R}^{n}\setminus E_{\boldsymbol{h}}$. In particular, $|\nabla F|,|F|$ are bounded near $E_{\boldsymbol{h}}$. We can extend $|F|$ to $E_{\boldsymbol{h}}$ continuously by setting $|F|=0$ on $E_{\boldsymbol{h}}$. It remains to show $F=f_{\boldsymbol{h}}$. Because $\mathrm{d}F$ and $\mathrm{d}f_{\boldsymbol{h}}$ have the same asymptotic behavior at infinity, we obtain